OpenStudy (anonymous):

sqr11x+3-2x=0 only the 11x+3 is squared

OpenStudy (whpalmer4):

Is that \[(11x+3)^2-2x=0\] or something else? The equation editor button really isn't hard to use...

OpenStudy (anonymous):

\[\sqrt{11x+3}-2x=0\] this is what i meant